Binomial probability:
Jessie is a fantastic goalie who stops three out of every four penalty shots taken on her. In a particularly tough game, she faced five penalty shots. What is the probability she lets only one of the five shots score?
If R=1 N=5, would probability of success be: P=3/4 and probability of failure be: Q=1/4 because I'm not sure 100%?

(I'm just having a hard time figuring out if I am writing my fractions correctly for: P=Probability of success and Q=Probability of failure. Therefore, I don't know if my equation should be set up as: 5C1(3/4)^1(1/4)^4 and if my answer than should be 1.46%. Please let me know if this is correct or not. Thank you and I really appreciate it too.)

Answer by stanbon(75887)

Binomial probability:
Jessie is a fantastic goalie who stops three out of every four penalty shots taken on her.
---
P(blocks shot) = 3/4 = 0.75
P(allows goal) = 0.25
----------
In a particularly tough game, she faced five penalty shots. What is the probability she lets only one of the five shots score?
---
Binomial Problem with n = 5 and p = 0.25 (allows goal)
P(allows one goal in 5 shots) = 5C1(0.25)(0.75)^4 = 0.3955
---
Comment: Get away from using the words "success" and "failure".
Use the words relating to the situation described. Actually
write the words down as I have above. That way you will aways
know what p is and what q is.
====================================
Cheers,
Stan H.